A biologist has two brine solutions, one containing 7% salt and another containing 28% salt. How many milliliters of each solution should she mix to obtain 1 L of a solution that contains 19.6% salt?
Answer by richwmiller(17219)

a+b=1,
0.07*a+0.28*b=0.196*1
a=1-b
0.07*(1-b)+0.28*b=0.196
0.07-0.07b+0.28*b=0.196
0.21*b=0.126
b=0.6
a=1-b
a=0.4 liter at 7% 400 mL
b=0.6 liter at 28% 600 mL
check
0.07*0.4+0.28*0.6=0.196*1
0.028+0.168=0.196
0.196=0.196
ok